Cost Behavior, Resource Usage, Excess Capacity

Rolertyme Company manufactures roller skates. With the exception of the rollers, all parts of the skates are produced internally. Neeta Booth, president of Rolertyme, has decided to make the rollers instead of buying them from external suppliers. The company needs 100,000 sets per year (currently it pays $1.90 per set of rollers).

The rollers can be produced using an available area within the plant. However, equipment for production of the rollers would need to be leased ($30,000 per year lease payment). Additionally, it would cost $0.50 per machine hour for power, oil, and other operating expenses. The equipment will provide 60,000 machine hours per year. Direct material costs will average $0.75 per set, and direct labor will average $0.25 per set. Since only one type of roller would be produced, no additional demands would be made on the setup activity. Other overhead activities (besides machining and setups), however, would be affected. The company’s cost management system provides the following information about the current status of the overhead activities that would be affected. (The supply and demand figures do not include the effect of roller production on these activities.) The lumpy quantity indicates how much capacity must be purchased should any expansion of activity supply be needed. The purchase price is the cost of acquiring the capacity represented by the lumpy quantity. This price also represents the cost of current spending on existing activity supply (for each block of activity).

Activity Price Cost Driver Supply Usage Lumpy
Quantity
Purchase
Purchasing Orders 25,000 23,000 5,000 $25,000
Inspection Hours 10,000 9,000 2,000 30,000
Materials handling Moves 4,500 4,300 500 15,000

Production of rollers would place the following demands on the overhead activities:

Activity Resource Demands
Machining 50,000 machine hours
Purchasing 2,000 purchase orders (associated with raw materials used to make the rollers)
Inspection 750 inspection hours
Materials handling 500 moves

Producing the rollers also means that the purchase of outside rollers will cease. Thus, purchase orders associated with the outside acquisition of rollers will drop by 5,000. Similarly, the moves for the handling of incoming orders will decrease by 200. The company has not inspected the rollers purchased from outside suppliers.

2. Calculate the total annual resource spending (for all activities except for setups) that the company will incur after production of the rollers begins. Break this cost into fixed and variable activity costs. In calculating these figures, assume that the company will spend no more than necessary.

Fixed Cost $
Variable Cost $
Total Cost $

What is the effect on resource spending caused by production of the rollers?

Decrease   = $

3. Refer to Requirement 2. For each activity, break down the cost of activity supplied into the cost of activity output and the cost of unused activity.

Activity Cost of Activity Supplied Cost of Activity Used Cost of Unused Activity
Machining $ $ $
Purchasing
Inspection
Materials handling
Answer #1

ANSWER

2).

Total annual resources spending
Activity * Fixed costs ** Variable costs Total costs
material usage - 75000 75000
labor - 25000 25000
machining 30000 25000 55000
purchasing 100000 - 100000
inspection 150000 - 150000
material handling 150000 - 150000
Totals $430000 $125000 $555000
* machining : lease payment of $30000

** material usage= 0.75 * 100000 = $75000

labor = 0.25 * 100000 =25000

machining= 0.50 * 50000 = $25000

purchasing : 23000 - 5000 + 2000 = 20000 the required supply. Since the resource is purchsed in units of 5000, the supply can be reduced from 25000 to 20000 . Cost of each block of resources is $25000. so, resource spending is (20000/5000)*25000=$100000.

Inspection : 9000 +750 = 9750 as the demand. Since the resource is purchased in blocks of 2000, the supply should be equal 10000. Thus, resource spending is (10000/2000)+$30000=#150000.

Material handling : demand = 4300 + 500 - 200 = 4600. Since the resource is purchased in blocks of 500, the required supply should be 5000. Thus, resource spending is (5000/500)*$15000=$150000.

Effect on resource spending of decision to produce rollers :
material (75000)
labor (25000)
machining (55000)
purchasing (effect is a saving) 25000 *
inspection 0 **
material handling (15000) ***
outside purchase (effect is a saving) 190000 ****
Total decrease $45000

* supply drops from 25000 to 20000 orders, saving $25000.

** Activity stays at 10000 hrs, no change in spending is needed.

*** Activity increases by 500 moves, spending increases by $15000.

**** Resource spending for outside purchase vanishes, saving $190000 ($1.90*100000).

3).

Activity Cost of activity supplied Cost of activity used Cost of unused activity*
machining $30000 $25000 $5000
purchasing 100000 100000 0
inspection 150000 146250 3750
material handling 150000 138000 12000

* Multiply the fixed activity rate by unused capacity :

          machining: ($30000/60000)*10000

          purchasing : ($100000/20000)*0

          inspection : ($150000/10000)*250

          material handling : ($150000/5000)*400

Note : The cost of activity used is computed by multiplying the fixed activity rate by the amount used or by subtracting the unused activity cost of activity supplied.

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************